OpenStudy (anonymous):

Which is not deposited by wind? A)loess B) longitudinal dunes C)till D)barchan dunes

OpenStudy (anonymous):

@tylermcmullen23

OpenStudy (anonymous):

which 2 do you KNOW that CAN be deposited by wind??

OpenStudy (anonymous):

the dunes

OpenStudy (anonymous):

well, i know that A can be deposited by wind. its an accumulation caused by wind depositation.. what the heck is a till?? lol

OpenStudy (anonymous):

couldn't tell ya lol

OpenStudy (anonymous):

oh lord lol. hold on ill find out lol

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

I think it's longitudinal waves maybe idk lol

OpenStudy (anonymous):

no. what i found is that tills are deposited directly from glacial ice

OpenStudy (anonymous):

so itd be till because its deposited from glacial ice, not wind

OpenStudy (anonymous):

okay thank you!!

OpenStudy (anonymous):

anytime! :)

OpenStudy (anonymous):

I have 1 more!

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

oops idk what just happened

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Why are desert landscapes jagged and rocky? The evaporation process creates angular surfaces. Lack of vegetation causes sinkholes between rock formations. There is little chemical weathering to round the landscape

OpenStudy (anonymous):

sinkholes are caused by some sort of collapse in the earth layers, not lack of vegetation, so b is incorrect. and evaporation wouldnt cause angular surfaces, bc it just makes the surface dry. so id say C

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 2x - 6. Find (g o f) (1) A) 0 B) 3 C) 23 D)–4

OpenStudy (anonymous):

composition of functions??? is that what its called?

OpenStudy (anonymous):

I think so yes

OpenStudy (anonymous):

not too bad.. i think i know this

OpenStudy (anonymous):

sweet!!

OpenStudy (anonymous):

im not 100% sure but try 23

OpenStudy (anonymous):

kk thanks! its a little test so it wont tell me if I am right or not until the end

OpenStudy (anonymous):

If the parent function is the square root of x, describe the translation of the function square root of x-1. One unit up from square root of x One unit to the right of square root of x One unit to the left of square root of x One unit down from square root of x

OpenStudy (anonymous):

since its x, it cant be down or up, bc that would be y. in order for it to be to the left, it would have to be sqrt(x+1) but since its sqrt(x-1), you would be to the right

OpenStudy (anonymous):

so i believe its B

OpenStudy (anonymous):

okay thank you:)

OpenStudy (anonymous):

youre welcome! anymore?

OpenStudy (anonymous):

How is the graph of y=square root of x=3 translated from the graph of y=square root of x shifted 3 units left shifted 3 units up shifted 3 units down shifted 3 units right

OpenStudy (anonymous):

is that"=" supposed to be a +??? or no?

OpenStudy (anonymous):

no they are equal signs

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

well, since it want "y=", it cant be moved left or right. so itd have to be between B and C.

OpenStudy (anonymous):

im not for sure on this one as far as up or down.. but i believe it is 3 units up

OpenStudy (anonymous):

What is the domain and range for the following function and its inverse? f(x)=3x-1/2

OpenStudy (anonymous):

options or no?

OpenStudy (anonymous):

yes but they are so long to type out lol

OpenStudy (anonymous):

domain: \[-\infty<x<\infty\] range: \[-\infty<f(x)<\infty\] inverse: \[\frac{ 2x+1 }{6}\] if those are options put together, idk

OpenStudy (anonymous):

a) f(x)=domain: all real numbers, range: all real numbers f^-1=domain: all real numbers, range: all real numbers b) f(x)=domain: x<0 range:y>-1 c) f(x) domain:x>0, range: y>0 d) f(x) domain: x>0, range: y<-1 domain: x<-1 range:y>0

OpenStudy (anonymous):

itd have to be A, because the domain and range that i got were both the same, which is \[\infty\] also known as all real numbers so.. i guess itd be a

OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 6. Find (f o g) (-5) 3 –59 26 –6

OpenStudy (anonymous):

well, the result of it is 8x-19. so plug in -5, whichmakes the equation now 8(-5)-19. 8*-5=-40. -40-19=-59

OpenStudy (anonymous):

u da bomb

OpenStudy (anonymous):

doubt it lol

OpenStudy (anonymous):

rewrite y=the square root of 9x+45-2 to make it easy to graph using a translation. Describe the graph.

OpenStudy (anonymous):

yeahh.. about that one.. lol idk does it look like \[y=\sqrt{(9x+45)-2}\] or \[y=\sqrt{9x+45-2} \]

OpenStudy (anonymous):

the options are like \[y=3\sqrt{x+5}-2\]

OpenStudy (anonymous):

im sorry meg, i dont know this one :(

OpenStudy (anonymous):

just pick c. its almost always right lol

OpenStudy (anonymous):

What is the domain and range for the following function and its inverse? f(x) = x2 – 2

OpenStudy (anonymous):

thast supposed to be x^2

OpenStudy (anonymous):

domain: all real numbers range: \[y \ge-2\]

OpenStudy (anonymous):

I hate math so much

OpenStudy (anonymous):

i do too. are those an option for one?

OpenStudy (anonymous):

Rewrite y = \[\sqrt{25x-75}\] + 3 to make it easy to graph using a translation. Describe the graph.

OpenStudy (anonymous):

ill try. because i think youre supposed to just simplify it. so ill try ok?

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

is this an option? \[5*\sqrt{x-3+3}\]

OpenStudy (anonymous):

yes but you have to say which way it translates too

OpenStudy (anonymous):

im not sure where it translates to